[Brahms-l] Shift report 20040417 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sat Apr 17 2004 - 07:49:41 EDT
------------------------------------------------------------
Date: Sat Apr 17 07:49:41 2004, Shift: 00:00-08:00
Supervisor: hiro, with: 
Log book#15 pages 107 - 107

Shift Summary:
------------------------------------------------------------
1.  no beam.
    Beam was stored around 5AM and spposed steered for our IR But our rate was about 1/20th~40th of STAR and Phenx rate. (about 80-90Hz) This looks like the situation in the last year.  
2.  no
3.  HVs for INL counter trips all the time.  The return values are not the same as the set values (at all).  We need to replace the mainframe and/or some pods (modules).


------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Apr 17 07:49:53 2004

This archive was generated by hypermail 2.1.8 : Sat Apr 17 2004 - 07:50:16 EDT